OpenStudy (anonymous):

18 is what percent of 40? 45 is 15% of what number 16% of what number is 20 what is 57% of 240

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

I think the first one is 45%

OpenStudy (mathstudent55):

Let's deal with "18 is what percent of 40?" first

OpenStudy (anonymous):

do you know the is over of method?

OpenStudy (mathstudent55):

Percent is a ratio. The entire amount is 100% That is 45. You want to know what percent is 18 out of 45. Set up a proportion: \(\dfrac{18}{45} = \dfrac{x}{100} \)

OpenStudy (anonymous):

cross multiply

OpenStudy (anonymous):

18 x 100 is?

OpenStudy (mathstudent55):

Now you can solve the proportion by cross multipying: \(45x = 18 \cdot 100\) \(45x = 1800\) x = 40 Answer: 40%

OpenStudy (anonymous):

no 40

OpenStudy (anonymous):

not 45

OpenStudy (mathstudent55):

Wait, I just realize I used 45 instead of 40. Let me do it again.

OpenStudy (anonymous):

so write this down: 18/40 = x/100

OpenStudy (mathstudent55):

Percent is a ratio. The entire amount is 100%; that is 40. You want to know what percent is 18 out of 40. Set up a proportion: \(\dfrac{18}{40} = \dfrac{x}{100} \)

OpenStudy (anonymous):



OpenStudy (anonymous):

It equals 45%

OpenStudy (mathstudent55):

40x = 18*100 40x = 1800 x = 45 Answer: 45% @mjperez18 You were correct.

OpenStudy (anonymous):

what is 18 x 100

OpenStudy (mathstudent55):

Ready for the second one?

OpenStudy (anonymous):

Okay on to the next one

OpenStudy (mathstudent55):

45 is 15% of what number

OpenStudy (anonymous):

the next one is a little bit different but still can use a proportion

OpenStudy (anonymous):

45 is 15% of what number

OpenStudy (mathstudent55):

Now 45 is the part, and is also15%. The entire number is x, unknown, and is also 100%.

OpenStudy (anonymous):

45 = 15 x = 100

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

now cross multiply

OpenStudy (anonymous):

45 x 100 =?

OpenStudy (mathstudent55):

We set up a new proportion. \(\dfrac{45}{x} = \dfrac{15}{100} \)

OpenStudy (anonymous):

3? or something to do with the number 3

OpenStudy (anonymous):

lets just start with 45 x 100

OpenStudy (anonymous):

300

OpenStudy (anonymous):

45 x 100 = 4500

OpenStudy (mathstudent55):

Again, cross multiply and solve for x. 15x = 45 * 100 15x = 4500 x = 300

OpenStudy (anonymous):

15 times x = 15x

OpenStudy (mathstudent55):

Correct. Answer is 300.

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

okay 16%of what number is 20

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

The third problem is just like the second one.

OpenStudy (anonymous):

okay so 20/x 16/100

OpenStudy (anonymous):

remember this ( is/ of = x/100

OpenStudy (mathstudent55):

16% of what number is 20 The unknown is the entire number, or 100%. The known is that 15% is the part and is 20

OpenStudy (mathstudent55):

\(\dfrac{20}{x} = \dfrac{16}{100} \)

OpenStudy (mathstudent55):

16x = 20 * 100 16x = 2000 x = 125

OpenStudy (anonymous):

125?

OpenStudy (mathstudent55):

@mjperez18 Your proportion and answer are correct.

OpenStudy (anonymous):

and the next is like the other 2

OpenStudy (mathstudent55):

Last one: "what is 57% of 240" This problem is different from all other three above, but it's the easiest percent problem there is.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

To find a percent of a number, just multiply the percent (written as a decimal) by the number. 57% is the same as 0.57 57% of 24 is the same as 0.57 * 240 All you need to do is that simple multiplication.

OpenStudy (mathstudent55):

^i meant 240 above not 24.

OpenStudy (anonymous):

13.68

OpenStudy (anonymous):

percent

OpenStudy (mathstudent55):

Correct. You did great!

OpenStudy (mathstudent55):

Do you still want to work on the problem of turning $750 into $1000 in 8 years, and how much the simple interest rate must be?

OpenStudy (anonymous):

yes please

OpenStudy (mathstudent55):

Ok. Can you please start a new post with just that question. I'll look for it, and I'll go there and will help you.

OpenStudy (anonymous):

I thought it was 950

OpenStudy (mathstudent55):

950? The question is asking for an interest rate in percent.

OpenStudy (anonymous):

well I have limited time so can we just complete the question here